# Efficient Approximate Degenerate Ordered Statistics Decoding for Quantum Codes via Reliable Subset Reduction

### v1

@@ III-A Reliability order based on BP4
Suppose that BP fails to provide a valid error after a maximum of $T$ iterations. At this point, we have the output distribution ${\mathbf{q}}_{i}=(q_{i}^{I},q_{i}^{X},q_{i}^{Y},q_{i}^{Z})$, $i=1,2,\dots,n$ from the last BP iteration. Additionally, we have the hard-decision results for all $T$ iterations, denoted by ${\mathbf{h}}_{i}^{(j)}\in\{I,X,Y,Z\}$ for $i=1,\dots,n$ and $j=1,\dots,T$. For example, ${\mathbf{h}}_{i}^{(1)},{\mathbf{h}}_{i}^{(2)},\dots,{\mathbf{h}}_{i}^{(T)}$ represent the hard-decision result history of BP at qubit $i$ and ${\mathbf{h}}_{1}^{(j)},{\mathbf{h}}_{2}^{(j)},\dots,{\mathbf{h}}_{n}^{(j)}$ represent the hard-decision results of BP at the $j$-th iteration.
Our objective is to assign reliability for each Pauli $X$ and $Z$ error. Following the reliability definitions in our previous paper [1], we define two types of measures in the following.
@@
In other words, $\bm{\eta}_{i}$ is the length of the last run in the string ${\mathbf{h}}_{i}^{(1)},{\mathbf{h}}_{i}^{(2)},\dots,{\mathbf{h}}_{i}^{(T)}$. For example, if the hard-decision outputs at the first qubit over the last five iterations are $X,Y,Z,Z,Z$, then the final hard decision output is $Z$ and $\bm{\eta}_{1}=3$. This means that the error at qubit 1 persisted unchanged as $Z$ for three iterations when BP stops.
For implementation, storing the entire hard-decision history ${\mathbf{h}}$ is unnecessary. One can record only the hard-decision outputs at the latest iteration and uses the latest belief vector ${\mathbf{q}}_{i}$ to update the values of $\{\bm{\eta}_{i}:i=1,\dots,n\}$.
Our modified BP4 algorithm, which incorporates hard-decision reliability measures, is presented in Algorithm 1.
Algorithm 1 BP4
Input: check matrix ${\mathbf{S}}\in\{0,1\}^{m\times 2n}$, syndrome ${\mathbf{z}}\in\{0,1\}^{m\times 1}$, depolarizing error rate $\epsilon$, maximum number of iterations $T$.
Output: A valid error estimate $\hat{{\mathbf{E}}}\in\{0,1\}^{2n}$ and belief distributions $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$.
Initialization:
Steps:
for $j=1$ to $T$ do
     Update $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$, using ${\mathbf{S}},{\mathbf{z}},$ and $\{{\mathbf{p}}_{i}\}_{i=1}^{n}$.
     for $i=1$ to $n$ do
         end if
     end for
     if  the hard-decision ${{\mathbf{h}}}$ matches the syndrome ${\mathbf{z}}$ then
         return $\varphi({\mathbf{h}})$ and $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$.
     end if
end for
return $\mathbf{0}$ and $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$. $\triangleright$ BP fails.
When BP fails, we are left with belief distributions ${\mathbf{q}}_{i}=(q_{i}^{I},q_{i}^{X},q_{i}^{Y},q_{i}^{Z})$ for $i=1,2,\dots,n$, which will be utilized in the reliability measure.
@@
Let ${\mathbf{E}}$ be the hard-decision output from BP4 after $T$ iterations. If ${\mathbf{E}}$ does not match the syndrome, OSD will be applied. We now present our OSD algorithm based on the reliability orders as in Algorithm 2, which is also referred to as OSD4-00. If the bits in the reliable subset are correct, the output of OSD4-00 is the required solution.
@@
Input: check matrix ${\mathbf{S}}\in\{0,1\}^{m\times 2n}$ of rank $n-k$, syndrome ${\mathbf{z}}\in\{0,1\}^{m\times 1}$, hard-decision reliability vector $\bm{\eta}\in\mathbb{Z}^{n}$, hard-decision vector ${\mathbf{E}}\in\{0,1\}^{2n}$, and belief distributions $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$.
@@
1:Sort the error bits in ascending order of reliability. Construct a column permutation function $\pi$ corresponding to the sorting order. Calculate $\pi({\mathbf{E}})$ and $\pi({\mathbf{S}}\Lambda)$.
2:Perform Gaussian elimination on $\pi({\mathbf{S}}\Lambda)$ and let the output be denoted by $R(\pi({\mathbf{S}}\Lambda))$, where $R$ is the corresponding row operations in the Gaussian elimination. If necessary, apply a column permutation function $\mu$ to ensure that the first $n-k$ columns of the resulting matrix are linearly independent. Thus we have
$\mu(R(\pi({\mathbf{S}}\Lambda)))=\left[\begin{array}[]{cc}{\mathbf{I}}_{n-k}&{\mathbf{A}}\\ \mathbf{0}&\mathbf{0}\end{array}\right],$ (5)
@@
3:Let ${\mathbf{z}}^{\prime}=R({\mathbf{z}})$ so that the parity-check condition (2) becomes ${\mathbf{z}}^{\prime}=\mu(R(\pi({\mathbf{S}}\Lambda)))(\mu(\pi(\hat{{\mathbf{E}}})))^{\top}$. Remove the last $m-(n-k)$ entries of ${\mathbf{z}}^{\prime}$.
4:Suppose $\mu(\pi({\mathbf{E}}))=\left[{\mathbf{E}}^{U}~{}~{}{\mathbf{E}}^{R}\right]$, where ${\mathbf{E}}^{U}$ represents the independent and unreliable subset of $n-k$ bits and ${\mathbf{E}}^{R}$ is the reliable subset of $n+k$ bits. return
$\hat{{\mathbf{E}}}\triangleq\pi^{-1}\left(\mu^{-1}\left(\left[{{\mathbf{z}}^{\prime}}^{\top}\oplus{{\mathbf{E}}^{R}}{\mathbf{A}}^{\top}\qquad{\mathbf{E}}^{R}\right]\right)\right).$ (6)
For comparison with conventional OSD procedures that do not incorporate the history of BP hard-decision outputs across iterations, we also define a reliability order that accounts for only the soft reliability functions of BP4.
@@
An OSD algorithm that follows a reliability order depending only on the belief distributions from the last BP iteration is referred to as mOSD. The OSD based on Definition 5 is referred to as mOSD4.
Note that the reliability metric in Definition 5 is distinct from those in [42, 43]. Other alternative metrics can also be considered, such as
@@ III-B OSD4-$w$
Given an OSD4-0 output vector, flipping one of its reliable bits to generate an error candidate takes time ${\cal O}(n)$. Similarly, generating an error candidate corresponding to a child node from its parent node also takes ${\cal O}(n)$ time.
@@
Consider the vector $\Big{[}{\mathbf{z}}^{\prime\top}\oplus{\mathbf{E}}^{R}{\mathbf{A}}^{\top}\,\Big{|}{\mathbf{E}}_{1}^{R}\,{\mathbf{E}}_{2}^{R}\,\cdots\,{\mathbf{E}}_{n+k}^{R}\Big{]}$ generated by $\operatorname{OSD}_{4}$-0 in (6), where ${\mathbf{E}}_{j}^{R}$ are reliable bits. Suppose that the matrix ${\mathbf{A}}$ in (5) has columns ${\mathbf{A}}_{1},{\mathbf{A}}_{2},\dots,{\mathbf{A}}_{n+k}\in\{0,1\}^{n-k}$. Then flipping the bit ${\mathbf{E}}_{i}^{R}$ generates the vector
$\left[{\mathbf{z}}^{\prime\top}\oplus{\mathbf{E}}^{R}{\mathbf{A}}^{\top}\oplus{\mathbf{A}}_{i}^{\top}\middle|{\mathbf{E}}_{1}^{R}\,{\mathbf{E}}_{2}^{R}\,\cdots\Big{(}1\oplus{\mathbf{E}}_{i}^{R}\Big{)}\cdots\,{\mathbf{E}}_{n+k}^{R}\right],$ (9)
which is a valid error candidate after permutations. It can be checked that this calculation takes time $\mathcal{O}(n)$. ∎
@@
We prioritize flipping the least reliable bits first. If an error with a lower Hamming weight is encountered, we update the solution accordingly. However, if the error has the same Hamming weight as the current one, we retain the previous solution since the bit flipped in the previous step is deemed less reliable.
@@ III-C Complexity of OSD4
We analyze the complexity of OSD based on 00-ADD in the following. The complexity of OSD4-0 is dominated by the Gaussian elimination step, which is then ${\cal O}(2n\times m^{2})={\cal O}(n^{3})$, assuming $m={\cal O}(n)$. For OSD4-$w$ with $w>0$, the total number of possible flips is $\sum_{j=0}^{w}\binom{n+k}{j}$, which can be expressed as ${\cal O}((n+k)^{w})$ or simply ${\cal O}(n^{w})$.
In Lemma 6, when we flip a bit in the reliable part, it requires only $\mathcal{O}(n)$ calculations to find an error candidate. Additionally, by running through $\sum_{j=0}^{w}\binom{n+k}{j}$ possibilities using the DFS algorithm, the OSD4-$w$ algorithm has a complexity of $\mathcal{O}(n^{3}+n^{w+1})$, which simplifies to $\mathcal{O}(n^{3})$ when $w\leq 2$.
The above OSD complexity is only required when BP fails to converge. BP has a worst-case time complexity of $O(n\gamma T)$ using a sequential schedule, where $n$ is the number of error variables, $\gamma$ is the mean column weight of the check matrix, and $T$ is the maximum number of iterations. However, BP typically converges quickly, within $O(\log N)$ iterations, for sparse matrices. This is also the case for MBP4 as demonstrated in Table VI in the simulation section. Thus MBP4 has a nearly linear complexity in $n$ from simulations and has a high probability of convergence for many spare-graph quantum codes [12, 13].
@@
Input: check matrix ${\mathbf{S}}\in\{0,1\}^{m\times 2n}$ of rank $n-k$, syndrome ${\mathbf{z}}\in\{0,1\}^{m\times 1}$, hard-decision reliability vector $\bm{\eta}\in\mathbb{Z}^{n}$, hard-decision vector ${\mathbf{E}}\in\{0,1\}^{2n}$, belief distributions $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$, and soft reliability threshold $\theta$.
Output: shortened variables $\tilde{{\mathbf{z}}}$, $\tilde{{\mathbf{S}}}$, ${\mathbf{E}}^{H}$, and $\sigma$.
@@
3:Let $\tau$ be a row permutation such that $\tau(\sigma({\mathbf{S}}\Lambda))=\begin{bmatrix}\tilde{{\mathbf{S}}}&{\mathbf{B}}\\ \mathbf{0}&{\mathbf{C}}\\ \end{bmatrix}$ for $\tilde{{\mathbf{S}}}\in\{0,1\}^{m^{\prime}\times(2n-v)}$ with $m^{\prime}\leq m$. The $m-m^{\prime}$ rows $\begin{bmatrix}\mathbf{0}&{\mathbf{C}}\end{bmatrix}$ correspond to parity checks related only to the highly reliable bits. Assume that $\tau({\mathbf{z}})=\begin{bmatrix}{\mathbf{z}}^{\prime\prime}\\ {\mathbf{z}}^{H}\end{bmatrix}$, where ${\mathbf{z}}^{\prime\prime}\in\{0,1\}^{m^{\prime}\times 1}$, and ${\mathbf{z}}^{H}\in\{0,1\}^{(m-m^{\prime})\times 1}$ are the syndrome bits corresponding to $\begin{bmatrix}\mathbf{0}&{\mathbf{C}}\end{bmatrix}$.
@@
if ${\mathbf{z}}^{H}\neq{\mathbf{C}}\big{(}{\mathbf{E}}^{H}\big{)}^{\top}$ then return failure;
@@
$\tilde{{\mathbf{z}}}=$ $\tilde{{\mathbf{S}}}\big{(}\hat{{\mathbf{E}}}^{\prime}\big{)}^{\top},$ (10)
     where $\tilde{{\mathbf{z}}}={\mathbf{z}}^{\prime\prime}-{\mathbf{B}}\big{(}{\mathbf{E}}^{H}\big{)}^{\top}$.
     if (10) is solvable then return $\tilde{{\mathbf{z}}}$, $\tilde{{\mathbf{S}}}$, ${\mathbf{E}}^{H}$, and $\sigma$;
     else return failure.
@@
IV Efficient OSD with highly reliable subset reduction
As we increase the order $w$ of OSD, we can expect an improvement in the decoder’s performance. In particular, when $w=n+k$ and $\delta=n$, OSD4-$(n+k)$ using $n$-ADD becomes a maximum likelihood decoder. However, increasing the order $w$ also leads to an increase in time complexity. To address this issue, we propose an approach to reduce the problem size for OSD. Recall that OSD is used to solve a linear system with $2n$ binary variables and $n-k$ constraints. If a subset of binary variables is known, the problem size–and consequently the computational complexity–can be reduced. The number of binary variables in the reduced linear system is referred to as the effective length for OSD, which corresponds to the length of the remaining binary error variable vector.
We note that OSD is applied when BP fails to find a valid error that matches the syndrome. During the $T$ iterations of message passing in BP, some coordinates become relatively more reliable than others. We have the following definition of highly reliable coordinates.

### v2

@@ III-A Reliability order based on BP4
Suppose that BP fails to provide a valid error after a maximum of ${\mathrm{T}}$ iterations. At this point, we have the output distribution ${\mathbf{q}}_{i}=(q_{i}^{I},q_{i}^{X},q_{i}^{Y},q_{i}^{Z})$, $i=1,2,\dots,n$ from the last BP iteration. Additionally, we have the hard-decision results for all ${\mathrm{T}}$ iterations, denoted by ${\mathbf{h}}_{i}^{(j)}\in\{I,X,Y,Z\}$ for $i=1,\dots,n$ and $j=1,\dots,{\mathrm{T}}$. For example, ${\mathbf{h}}_{i}^{(1)},{\mathbf{h}}_{i}^{(2)},\dots,{\mathbf{h}}_{i}^{({\mathrm{T}})}$ represent the hard-decision result history of BP at qubit $i$ and ${\mathbf{h}}_{1}^{(j)},{\mathbf{h}}_{2}^{(j)},\dots,{\mathbf{h}}_{n}^{(j)}$ represent the hard-decision results of BP at the $j$-th iteration.
Our objective is to assign reliability for each Pauli $X$ and $Z$ error. Following the reliability definitions in [1, 43], we define two types of measures in the following.
@@
In other words, $\bm{\eta}_{i}$ is the length of the last run in the string ${\mathbf{h}}_{i}^{(1)},{\mathbf{h}}_{i}^{(2)},\dots,{\mathbf{h}}_{i}^{({\mathrm{T}})}$. For example, if the hard-decision outputs at the first qubit over the last five iterations are $X,Y,Z,Z,Z$, then the final hard decision output is $Z$ and $\bm{\eta}_{1}=3$. This means that the error at qubit 1 persisted unchanged as $Z$ for three iterations when BP stops.
When BP fails, the belief distributions ${\mathbf{q}}_{i}=(q_{i}^{I},q_{i}^{X},q_{i}^{Y},q_{i}^{Z})$ will be utilized in the reliability measure.
@@
Our modified BP4 algorithm is presented in Algorithm 1.
Algorithm 1 BP4
Input: check matrix ${\mathbf{H}}\in\{0,1\}^{m\times 2n}$, syndrome ${\mathbf{s}}\in\{0,1\}^{m\times 1}$, error rate $\epsilon$, maximum number of iterations ${\mathrm{T}}$.
Output: A valid error estimate $\hat{{\mathbf{e}}}\in\{0,1\}^{2n}$, belief distributions $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$ and hard-decision reliability vector $\bm{\eta}\in\mathbb{Z}^{n}$.
Initialization:
Steps:
for $j=1$ to ${\mathrm{T}}$ do
     Update $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$, using ${\mathbf{H}},{\mathbf{s}},$ and $\{{\mathbf{p}}_{i}\}_{i=1}^{n}$.
     for $i=1$ to $n$ do
         end if
     end for
     if  the hard-decision ${{\mathbf{h}}}$ matches the syndrome ${\mathbf{s}}$ then
         return “Success" and $\varphi({\mathbf{h}})$.
     end if
end for
return “Failure", $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$, and $\bm{\eta}$. $\triangleright$ BP fails.
@@
Let $\tilde{{\mathbf{e}}}$ be the hard-decision output from BP4 after ${\mathrm{T}}$ iterations. If $\tilde{{\mathbf{e}}}$ does not match the syndrome, OSD will be applied. We now present our OSD algorithm based on the reliability orders as in Algorithm 2, which is referred to as OSD4-00.
@@
Input: check matrix ${\mathbf{H}}\in\{0,1\}^{m\times 2n}$ of rank $n-k$, syndrome ${\mathbf{s}}\in\{0,1\}^{m\times 1}$, hard-decision reliability vector $\bm{\eta}\in\mathbb{Z}^{n}$, hard-decision vector $\tilde{{\mathbf{e}}}\in\{0,1\}^{2n}$, and belief distributions $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$.
@@
1:Sort the error bits in ascending order of reliability. Construct a column permutation function $\pi$ corresponding to the sorting order. Calculate $\pi(\tilde{{\mathbf{e}}})$ and $\pi({\mathbf{H}}\Lambda)$.
2:Perform Gaussian elimination on $\pi({\mathbf{H}}\Lambda)$ and let the output be denoted by $R(\pi({\mathbf{H}}\Lambda))$, where $R$ represents the corresponding row operations in the Gaussian elimination. If necessary, apply a column permutation function $\mu$ to ensure that the first $n-k$ columns of the resulting matrix are linearly independent. Thus we have
$\mu(R(\pi({\mathbf{H}}\Lambda)))=\left[\begin{array}[]{cc}{\mathbf{I}}_{n-k}&{\mathbf{A}}\\ \mathbf{0}&\mathbf{0}\end{array}\right],$ (7)
@@
3:Let ${\mathbf{s}}^{\prime}=R({\mathbf{s}})$ so that the parity-check condition (2) becomes
${\mathbf{s}}^{\prime}=\mu(R(\pi({\mathbf{H}}\Lambda)))(\mu(\pi({\mathbf{e}})))^{\top}.$ (8)
Remove the last $m-(n-k)$ entries of ${\mathbf{s}}^{\prime}$.
4:Let $\mu(\pi(\tilde{{\mathbf{e}}}))=\left[\tilde{{\mathbf{e}}}^{U}~{}~{}\tilde{{\mathbf{e}}}^{R}\right]$, where $\tilde{{\mathbf{e}}}^{U}$ represents the independent and unreliable subset of $n-k$ bits and $\tilde{{\mathbf{e}}}^{R}$ is the reliable subset of $n+k$ bits. Use the reliable subset $\tilde{{\mathbf{e}}}^{R}$ and the modified parity-check condition 8 to select feasible unreliable bits
$\tilde{{\mathbf{e}}}^{U}={{\mathbf{s}}^{\prime}}^{\top}\oplus{\tilde{{\mathbf{e}}}^{R}}{\mathbf{A}}^{\top}.$ (9)
return
$\hat{{\mathbf{e}}}\triangleq\pi^{-1}\left(\mu^{-1}\left(\left[{{\mathbf{s}}^{\prime}}^{\top}\oplus{\tilde{{\mathbf{e}}}^{R}}{\mathbf{A}}^{\top}\qquad\tilde{{\mathbf{e}}}^{R}\right]\right)\right).$ (10)
We also define a reliability order that accounts for only the soft reliability functions of BP4.
@@
The OSD based on Definition 5 is referred to as mOSD4.
Other alternative metrics can also be considered, such as
@@ III-B OSD4-$w$
Given an OSD4-0 output vector, flipping one of its reliable bits to generate an error candidate takes time $O(n)$. Similarly, generating an error candidate corresponding to a child node from its parent node also takes $O(n)$ time.
@@
Consider the vector $\Big{[}{\mathbf{s}}^{\prime\top}\oplus{\mathbf{e}}^{R}{\mathbf{A}}^{\top}\,\Big{|}\tilde{{\mathbf{e}}}_{1}^{R}\,\tilde{{\mathbf{e}}}_{2}^{R}\,\cdots\,\tilde{{\mathbf{e}}}_{n+k}^{R}\Big{]}$ generated by $\operatorname{OSD}_{4}$-0 in (10), where $\tilde{{\mathbf{e}}}_{j}^{R}$ are reliable bits. Suppose that the matrix ${\mathbf{A}}$ in (7) has columns ${\mathbf{A}}_{1},{\mathbf{A}}_{2},\dots,{\mathbf{A}}_{n+k}\in\{0,1\}^{n-k}$. Then according to (9), flipping the bit $\tilde{{\mathbf{e}}}_{i}^{R}$ corresponds to adding the OSD4-0 output by the vector
$\begin{bmatrix}{\mathbf{A}}_{j}^{\top}&0\cdots 0&1&0\cdots 0\end{bmatrix}\in\{0,1\}^{2n},$ (13)
where ${\mathbf{A}}_{j}$ is the $j$-th column of ${\mathbf{A}}$, and generating the vector
$\left[{\mathbf{s}}^{\prime\top}\oplus\tilde{{\mathbf{e}}}^{R}{\mathbf{A}}^{\top}\oplus{\mathbf{A}}_{i}^{\top}\middle|\tilde{{\mathbf{e}}}_{1}^{R}\,\tilde{{\mathbf{e}}}_{2}^{R}\,\cdots\Big{(}1\oplus\tilde{{\mathbf{e}}}_{i}^{R}\Big{)}\cdots\,\tilde{{\mathbf{e}}}_{n+k}^{R}\right],$ (14)
which is a valid error candidate after permutations. This calculation takes $O(n)$ time. ∎
@@ III-C Complexity of OSD4
We analyze the complexity of OSD based on 00-ADD in the following. The complexity of OSD4-0 is dominated by the Gaussian elimination step, which is then $O(2nm^{2})=O(n^{3})$, assuming $m=O(n)$. For OSD4-$w$ with $w>0$, the total number of possible flips is $\sum_{j=0}^{w}\binom{n+k}{j}$, which can be expressed as $O((n+k)^{w})$ or simply $O(n^{w})$.
In Lemma 6, when we flip a bit in the reliable part, it requires only $O(n)$ calculations to find an error candidate. Additionally, by running through $\sum_{j=0}^{w}\binom{n+k}{j}$ possibilities using the DFS algorithm, the OSD4-$w$ algorithm has a complexity of $O(n^{3}+n^{w+1})$, which simplifies to $O(n^{3})$ when $w\leq 2$.
@@
Input: check matrix ${\mathbf{H}}\in\{0,1\}^{m\times 2n}$ of rank $n-k$, syndrome ${\mathbf{s}}\in\{0,1\}^{m\times 1}$, hard-decision reliability vector $\bm{\eta}\in\mathbb{Z}^{n}$, hard-decision vector ${\mathbf{e}}\in\{0,1\}^{2n}$, belief distributions $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$, and soft reliability threshold $\theta$.
Output: shortened variables $\tilde{{\mathbf{s}}}$, $\tilde{{\mathbf{H}}}$, ${\mathbf{e}}^{H}$, and $\sigma$.
@@
3:Let $\tau$ be a row permutation such that $\tau(\sigma({\mathbf{H}}\Lambda))=\begin{bmatrix}\tilde{{\mathbf{H}}}&{\mathbf{B}}\\ \mathbf{0}&{\mathbf{C}}\\ \end{bmatrix}$ for $\tilde{{\mathbf{H}}}\in\{0,1\}^{m^{\prime}\times(2n-v)}$ with $m^{\prime}\leq m$. The $m-m^{\prime}$ rows $\begin{bmatrix}\mathbf{0}&{\mathbf{C}}\end{bmatrix}$ correspond to parity checks related only to the highly reliable bits. Assume that $\tau({\mathbf{s}})=\begin{bmatrix}{\mathbf{s}}^{\prime\prime}\\ {\mathbf{s}}^{H}\end{bmatrix}$, where ${\mathbf{s}}^{\prime\prime}\in\{0,1\}^{m^{\prime}\times 1}$, and ${\mathbf{s}}^{H}\in\{0,1\}^{(m-m^{\prime})\times 1}$ are the syndrome bits corresponding to $\begin{bmatrix}\mathbf{0}&{\mathbf{C}}\end{bmatrix}$.
@@
if ${\mathbf{s}}^{H}\neq{\mathbf{C}}\big{(}{\mathbf{e}}^{H}\big{)}^{\top}$ then return “Failure";
@@
$\tilde{{\mathbf{s}}}=$ $\tilde{{\mathbf{H}}}\big{(}\hat{{\mathbf{e}}}^{\prime}\big{)}^{\top},$ (15)
     where $\tilde{{\mathbf{s}}}={\mathbf{s}}^{\prime\prime}-{\mathbf{B}}\big{(}{\mathbf{e}}^{H}\big{)}^{\top}$.
     if (15) is solvable then return $\tilde{{\mathbf{s}}}$, $\tilde{{\mathbf{H}}}$, ${\mathbf{e}}^{H}$, and $\sigma$;
     else return “Failure".
@@
IV Highly reliable subset reduction
As we increase the order $w$ of OSD, we can expect an improvement in the decoder’s performance. In particular, when $w=n+k$ and $\delta=n$, OSD4-$(n+k)$ using $n$-ADD becomes a maximum likelihood decoder. However, increasing the order $w$ also leads to an increase in time complexity.
To address this issue, we propose an approach to reduce the problem size for OSD. Recall that OSD is used to solve a linear system with $2n$ binary variables and $n-k$ constraints. If a subset of binary variables is known, the problem size–and consequently the computational complexity–can be reduced. The number of binary variables in the reduced linear system is referred to as the effective length for OSD, which corresponds to the length of the remaining binary error variable vector. We have the following definition of highly reliable coordinates.